Two plates separated by a distance 16.3 mm are cha
Answer #1

Change in Potential = q deltaV

and deltaV = - 7.25 volts

deltaU = -8.30 x 10^-3 x -7.25 = 60.175 x 10^-3 J    ......Ans


Its potential energy increase.

And

Work done by force = F.d = 6.81 ( 16.3 x 10^-3 m) = 111 x 10^-3 J


Using work energy theorem,

Work done by force + work done by electric force = change in KE

(work done = - deltaU)

111 x 10^-3 + ( - 60.175 x 10^-3) = deltaKE

deltaKE Or deltaT = 50.83 J .......Ans

total enery change, deltaE = T + U = 111 x 10^-3 J .....Ans